ASSUME THAT THE LENGTH OF TIME . X. BETWEEN CHARGES OF A CELLULAR PHONE IS NORMALLY DISTIBUTED WITH A MEAN OF 10 HOURS AND A STANDARD DEVIATON OF 1.5 HOURS. FIND THE PROBABILITY THAT THE CELL PHONE WILL LAST BETWEEN 8 AND 12 HOURS BETWEEN CHARGES.
Answer by Boreal(15235)

z=(x-mean)/sd
=(8-10)/1.5=-1.33
=(12-10)/1.5=+1.33
probability z is between those two values is 0.8664